# Helix: Distributed Serving of Large Language Models via Max-Flow on Heterogeneous GPUs

### v1

@@ 3 Optimization Formulation in Helix
(a) A cluster of three nodes with model placement. Network connections between the coordinator and compute nodes transmit tokens (4 Byte) while other connections transmit intermediate activations (16 KB)
(b) Graph abstraction of the cluster.
Figure 2: Graph abstraction of a 3-node cluster with given model placement. Numbers on the edges in Fig. 2(b) represent their capacity, which is the number of tokens that can pass through the edges per second. Max flow between source and sink equals the max serving throughput of the cluster.
3.1 Formulation of LLM Serving
A cluster to serve LLMs generally contains one coordinator node $h$ and a group of compute nodes $\mathcal{C}$. Each compute node $c_{i}\in\mathcal{C}$ has a compute throughput $T_{i}$ and given GPU VRAM size. Network connections between nodes in the cluster have given their throughput and latency. Based on the cluster information, LLM serving requires finding a placement of model layers to compute nodes to maximize serving performance. The model placement function $\Psi:\mathcal{C}\mapsto\mathcal{P}(\mathcal{M})$ takes as input a compute node and returns a (usually continuous) subset of the model $\mathcal{M}$. One widely used metric [4, 31] to assess a model placement is the performance of the pipeline stage with the lowest compute capacity $\min_{i}\sum_{j}\text{capacity}(j)\cdot 1_{i\in\Psi(j)}$, which sums up the compute capacity of all instances serving one layer. As we will show below, only considering compute capacity yields sub-optimal model placements for heterogeneous clusters.
@@ 3.1 Formulation of LLM Serving
Before diving into our Max-Flow formulation of heterogeneous LLM serving, we first use an example to show why we need to co-optimize model partition, device placement, and request scheduling in a Max-Flow problem.
@@
In this example, the devices are shown in Fig. 1(a). There are two regions with a low bandwidth between them. Region 1 has a powerful A100 GPU, while Region 2 has a less powerful L4 GPU and three T4 GPUs, but has a high bandwidth within the region. The pairwise bandwidths are independent. If we, following the common approach, first partition the model, then assign devices to each partition, the placement plan will be as Fig. 1(b). In this plan, although the last pipeline stage has a T4 and an L4 GPU, its throughput is bound by the previous stage’s output throughput, which only has 2 T4 GPUs. This indicates a necessity to co-optimize the pipeline partition plan and placement of pipeline stages.
@@
However, even with a perfectly balanced compute capacity at each pipeline stage, as shown in Fig. 1(c), the solution can still be sub-optimal. In this solution, it assigns the powerful A100 to individually serve some layers, while other GPUs runs in a data parallel manner for the rest layers. However, communications from one pipeline stage to another become the bottleneck. For every request, its intermediate state is sent from Region 1 to Region 2 via low bandwidth. This eventually creates congestion on the A100’s send side. Instead, Fig. 1(d) assigns two T4 GPUs running in parallel with the A100. This splits the workload between the A100 GPU and the two T4 GPUs, so as to reduce the communication volume on the slow link.
3.2 Heterogeneous LLM Serving as Max-Flow
@@
To optimize model placement, Helix needs a way to determine the max serving throughput of different model placements. To achieve this, we transform a cluster of compute nodes with assigned model layers into a directed graph with edge capacity. The edge capacity denotes the number of tokens compute nodes and network connections can process and transmit per second. Max flow between source and sink vertices in the graph, which represent the coordinator node, gives us the max serving throughput of the cluster with current model placement. Below we show the formal construction of a cluster’s graph abstraction with a given model placement.
For a given cluster with coordinator node $h$, a set of compute nodes $\mathcal{C}$, and a model placement $\Psi$, we can transform entities in the cluster into elements of its graph abstraction as follows. An example of such a graph abstraction of a cluster with given model placement is shown in Fig. 2
Compute Nodes.
For each compute node $c_{i}\in\mathcal{C}$, we represent it with two connected vertices in the graph. We name the two vertices $c^{in}_{i}$ and $c^{out}_{i}$. The capacity of the directed edge $(c^{in}_{i},c^{out}_{i})$ represents the max number of tokens this node can process within one second. It is the minimum of the node’s compute and network throughput. Helix performs a one-time profiling to measure the throughput of all compute nodes.
Coordinator Node.
We represent the coordinator node as source and sink vertices in the graph.
Network Connection.
In a given cluster, a node may communicate with any other nodes, creating $O(|\mathcal{C}|^{2})$ possible directed network connections between different nodes. However, only a subset of those connections are valid based on the model placement as described below. A valid connection should satisfy one of the following three criteria: (1) the connection is from coordinator node $h$ to compute node $c_{i}$ and $c_{i}$ holds the first layer of the model; (2) the connection is from a compute node $c_{j}$ to coordinator node $h$ and $c_{j}$ holds the last layer of the model; (3) the connection is from one compute node $c_{i}$ to another compute node $c_{j}$ and $c_{j}$ holds model layers immediately needed after inference on $c_{i}$. For the first and second case, we represent the connection with directed edge $(source,c^{in}_{i})$ and $(c^{out}_{j},sink)$ respectively, with capacity equal to the connection bandwidth divided by the transmission size of a token (a few bytes). For the third case, we represent the connection with a directed edge $(c_{i}^{out},c_{j}^{in})$, and the capacity equals to the connection bandwidth divided by the transmission size of an activation (tens of kilobytes). This models the throughput constraint imposed by network connection speed between different nodes. We denote the full set of possible network connections as $\mathcal{E}$.
After constructing the equivalent graph abstraction of a cluster, we run preflow-push algorithm [6] to get the max flow between source and sink node. One unit of flow here represents one token that can pass through a compute node or network connection in one second. Therefore, the max flow gives us the max possible serving throughput of the cluster with current model placement.
Table 2: Variables used in MILP.
@@
Table 3: Constraints used in MILP.
@@
Infer. throughput $O(|\mathcal{C}|)$ $\sum_{u}f_{ui}\leq\sum_{j=1}^{k}b_{i}^{j}\cdot\text{throughput}_{j}$
@@
Trans. throughput $O(|\mathcal{E}|)$ $f_{i,j}\leq d_{i,j}\cdot\text{throughput}_{i,j}$
3.3 Optimal Model Placement with MILP
The previous section presented an approach for obtaining the max serving throughput of a cluster for a given model placement. In this section, we introduce a mixed-integer linear programming (MILP)-based method to find a model placement that maximizes the max flow, thus maximizing serving throughput. The MILP formulation has linear number of variables and constraints with respect to the number of compute nodes and network connections. Key challenges addressed include (1) formulation of system-level constraints as linear number of conditions to satisfy, (2) expression of these conditions with linear number of variables, and (3) linearization of each condition using auxiliary variables, specifically, each constraint is expressed as most three linear constraints with the help of at most two auxiliary variables. An overview of the variables and constraints is shown in Table. 2 and 3.
@@ 3.3 Optimal Model Placement with MILP
To represent the model placement on each compute node, we introduce two groups of variables in our MILP formulation. Suppose the model has a total of $L$ layers and each compute node holds a continuous subset of the model. For each compute node $c_{i}$, we introduce an integral variable $s_{i}$ to represent the first layer $c_{i}$ holds. Suppose compute node $c_{i}$ can hold at most $k$ layers on its GPU, we further introduce $k$ binary variables $b_{i}^{1},b_{i}^{2},...,b_{i}^{k}$ to indicate the number of layers node $c_{i}$ holds ($b_{i}^{j}=1$ if $c_{i}$ holds $j$ layers). We choose to express model placement with $k$ binary variables (instead of one integer for the number of layers) because this formulation facilitates the expression of inference throughput constraints as discussed below. The end layer index of $c_{i}$ can be expressed as $e_{i}=s_{i}+\sum_{j=1}^{k}j\cdot b_{i}^{j}$. Therefore, $c_{i}$ holds layers in range $[s_{i},e_{i})$.
Figure 3: Helix overview. In Helix, the coordinator plans model placement as described in Sec. 3.3. We only need to run model placement once for each cluster. When a new request arrives, the coordinator node runs Helix scheduler to assign it a per-request pipeline and sends the request to the first node in the pipeline. Each compute node in the pipeline performs inference on the request on the layers it is responsible for and sends the (output for the) request to the next node in the pipeline. When the last node in the pipeline finishes performing inference on its layers, it will send the output token for the request to the coordinator (Worker Finished). The coordinator schedules generation of the next token for the request using the same pipeline.
@@
We introduce two groups of variables to constrain the number of inference requests that can go through each network connection in the cluster. For network connection between compute node $c_{i}$ and $c_{j}$, we introduce a real variable $f_{i,j}$ to denote the amount of flow from $c_{i}^{out}$ to $c_{j}^{in}$ in the graph abstraction. We further introduce a binary variable $d_{i,j}$ to denote whether the network connection is valid (as defined in Sec. 3.2). The constraints we introduce below will use $d_{i,j}$ to ensure that requests can only be transmitted through valid connections. For network connections between coordinator node and compute nodes, we similarly introduce two variables as above, but replace $i$/$j$ with source/sink.
@@
To ensure that the model placement found by the MILP solver is valid, we need the following two constraints for each compute node $c_{i}$. First, $c_{i}$ should have only one valid model placement, meaning that $\sum_{j=1}^{k}b_{i}^{j}=1$. Moreover, the first and last layer $c_{i}$ holds must be within the range of $L$ layers, meaning that $0\leq s_{i}<L$ and $e_{i}\leq L$.
@@
For each compute node $c_{i}$, the sum of flow that goes in to $c_{i}^{in}$ must equal to that goes out of $c_{i}^{out}$ because of flow conservation. This constraint can be expressed as $\sum_{u}f_{u,i}=\sum_{v}f_{i,v}$, where $u$ and $v$ enumerates through all nodes except $i$.
@@
For compute node $c_{i}$, the amount of flow that passes through $(c_{i}^{in},c_{i}^{out})$ should be no larger than its maximum inference throughput. We can impose this constraint with $\sum_{u}f_{ui}\leq\sum_{j=1}^{k}b_{i}^{j}\cdot T_{j}$. Here, $T_{j}$ is a constant that represents the maximum number of tokens node $c_{i}$ can process in one second when holding $j$ layers, which is obtained through a one-time profiling process.
@@
We need to determine the validity of network connections to know whether requests can be transmitted through them. For a network connection from the coordinator node to compute node $c_{i}$, it is valid only if $c_{i}$ holds the first layer of the model. To express this constraint with MILP, we need to linearize it into the following form: $s_{i}\leq L\cdot(1-d_{source,i})$. Similarly, for network connection from compute node $c_{i}$ to coordinator, we constrain its validity with $L\cdot d_{i,sink}\leq e_{i}$. For network connection from compute node $c_{i}$ to $c_{j}$, its validity $d_{i,j}$ is determined by whether $s_{j}\leq e_{i}<e_{j}$ holds. To linearize this condition, we need to introduce two binary auxiliary variables $cond^{1}_{i,j}$ and $cond^{2}_{i,j}$. $cond^{1}_{i,j}$ takes value 1 only if $s_{j}\leq e_{i}$, which can be linearized as $(L+1)(1-cond^{1}_{i,j})\geq s_{j}-e_{i}$. $cond^{2}_{i,j}$ takes value 1 only if $e_{i}<e_{j}$, which can be linearized as $e_{j}-e_{i}\geq 1-(L+1)(1-cond^{2}_{i,j})$. The network connection is valid only if both binary auxiliary variables holds true, which can be expressed as $d_{i,j}\leq 0.5*cond^{1}_{i,j}+0.5*cond^{2}_{i,j}$.
We remark that if $s_{j}<e_{i}<e_{j}$, then requests coming from $c_{i}$ will only infer layers $[e_{i},e_{j})$ on $c_{j}$. We call this partial inference. If partial inference is not allowed, then the connection validity constraints can be simplified to $d_{i,j}=1$ only if $e_{i}=s_{j}$, which linearizes to two constraints $L\cdot d_{i,j}\leq L+s_{j}-e_{i}$ and $L\cdot d_{i,j}\geq L-s_{j}+e_{i}$.
@@
We only allow flow to pass through valid network connections and the flow that passes should not be larger than the connection’s maximum transmission throughput. To enforce this constraint, we add $f_{i,j}\leq d_{i,j}\cdot S_{i,j}$ as a constraint into the MILP problem. $S_{i,j}$ is the maximum number of tokens that can be transmitted through the network connection, which can be calculated via profiling and using methods mentioned in Sec. 3.2.
@@
3.4 Analyzing and Speeding up MILP

### v2

@@ 4. Optimization Formulation in Helix
4.1. Formulation of LLM Serving
A cluster to serve LLMs generally contains one coordinator node $h$ and a group of compute nodes $\mathcal{C}$. Each compute node $c_{i}\in\mathcal{C}$ has a compute capacity and GPU VRAM size. Compute nodes with multiple GPUs can be abstracted as a single logical node, aggregating GPUs’ combined computational capacity and GPU VRAM resources. Throughput and latency of network connections between nodes in the cluster are also given. Based on the cluster information, LLM serving requires finding a placement of model layers to compute nodes to maximize a serving performance metric that will be defined below. The model placement function $\Psi:\mathcal{C}\mapsto\mathcal{P}(\mathcal{M})$ takes as input a compute node and returns a (usually continuous) subset of the model $\mathcal{M}$. Here, we assume pipeline parallelism for inter-node parallelization and tensor parallelism for intra-node parallelization across GPUs, as tensor parallelism demands extensive communication and may be constrained by network. One widely used metric (borzunov2022petals, ; ryabinin2023swarm, ) to assess a model placement is the performance of the pipeline stage with the lowest compute capacity $\min_{i}\sum_{j}\text{capacity}(j)\cdot 1_{i\in\Psi(j)}$, where capacity$(j)$ is the compute capacity of the $j^{th}$ node. As we will show below, only considering compute capacity yields sub-optimal model placements for heterogeneous clusters.
@@ 4.1. Formulation of LLM Serving
4.2. Necessity of Joint Optimization
@@
Before diving into our Max-Flow formulation of heterogeneous LLM serving, we first use an example to show why we need to co-optimize model partition, device placement, and request scheduling as a Max-Flow problem. In this example, the clusters are shown in Fig. 1(a). There are two regions with a low bandwidth between them. Region 1 has a powerful A100 GPU, while Region 2 has a less powerful L4 GPU and three T4 GPUs, but has a high bandwidth within the region. The pairwise bandwidths are independent. If we follow the common approach, which statically partitions the model and then assigns devices to each partition, the placement plan will be as Fig. 1(b). In this plan, although the last pipeline stage has a T4 and an L4 GPU, its throughput is bound by the previous stage’s output throughput, which only has 2 T4 GPUs. This indicates a necessity to co-optimize the pipeline partition plan and placement of pipeline stages.
@@
However, even with a perfectly balanced compute capacity at each pipeline stage, as shown in Fig. 1(c), the solution can still be sub-optimal. In this solution, it assigns the powerful A100 to individually serve some layers, while other GPUs run in a data parallel manner for the rest of the layers. However, communications from one pipeline stage to another become a bottleneck. For every request, its intermediate state is sent from Region 1 to Region 2 via low bandwidth. This eventually creates congestion on the A100’s send side. Instead, Fig. 1(d) assigns two T4 GPUs running in parallel with the A100. This divides the workload between the A100 GPU and the two T4 GPUs, reducing communication on the slow link.
4.3. Heterogeneous LLM Serving as Max-Flow
@@
To optimize model placement, Helix needs a way to determine the max serving throughput of different model placements. To achieve this, we transform a cluster of compute nodes with assigned model layers into a directed graph with edge capacity. The edge capacity denotes the number of tokens compute nodes and network connections can process/transmit per second. Max flow between source and sink vertices in the graph, which represent the coordinator node, gives us the max serving throughput of the cluster with the current model placement. The following shows the formal construction.
Table 5. Variables used in MILP.
@@
For a given cluster with coordinator node $h$, a set of compute nodes $\mathcal{C}$, and a model placement $\Psi$, we can transform entities in the cluster into elements of its graph abstraction as follows. An example of such a graph abstraction of a cluster with given model placement is shown in Fig. 2(b)
Compute and coordinator nodes.
For each compute node $c_{i}\in\mathcal{C}$, we represent it with two connected vertices in the graph. We name the two vertices $c^{in}_{i}$ and $c^{out}_{i}$. The capacity of the directed edge $(c^{in}_{i},c^{out}_{i})$ represents the max number of tokens this node can process in one second. It is the minimum of the node’s compute and network throughput. Helix performs a one-time profiling to measure the throughput of all compute nodes. For the coordinator node, we represent it as source and sink vertices in the graph.
Network connections.
In a given cluster, a node may communicate with any other nodes, creating $O(|\mathcal{C}|^{2})$ possible directed network connections between different nodes. However, only a subset of those connections are valid based on the model placement as described below. A valid connection should satisfy one of the following three criteria: (1) the connection is from coordinator node $h$ to compute node $c_{i}$ and $c_{i}$ holds the first layer of the model; (2) the connection is from a compute node $c_{j}$ to coordinator node $h$ and $c_{j}$ holds the last layer of the model; (3) the connection is from one compute node $c_{i}$ to another compute node $c_{j}$ and $c_{j}$ holds model layers immediately needed after inference on $c_{i}$. For the first and second case, we represent the connection with directed edge $(source,c^{in}_{i})$ and $(c^{out}_{j},sink)$ respectively, with capacity equal to the connection bandwidth divided by the transmission size of a token (a few bytes). For the third case, we represent the connection with a directed edge $(c_{i}^{out},c_{j}^{in})$, and the capacity equals the connection bandwidth divided by the transmission size of an activation (tens of kilobytes). Helix performs a one-time profiling and uses the average bandwidth as the connection bandwidth. The capacity of the edges models the throughput constraint imposed by the speed of network connection between different nodes. We denote the full set of possible network connections by $\mathcal{E}$.
After constructing the equivalent graph abstraction of a cluster, we run the preflow-push algorithm (cheriyan1989analysis, ) to get the max flow between source and sink node. One unit of flow here represents one token that can pass through a compute node or network connection in one second. Therefore, the max flow gives us the max possible serving throughput of the cluster with current model placement.
Table 6. Constraints used in MILP.
@@
Infer. throughput $O(|\mathcal{C}|)$ $\sum_{u}f_{ui}\leq\sum_{j=1}^{k}b_{i}^{j}\cdot T_{j}$
@@
Trans. throughput $O(|\mathcal{E}|)$ $f_{i,j}\leq d_{i,j}\cdot S_{i,j}$
4.4. Optimal Model Placement with MILP
The previous section presented an approach for obtaining the max serving throughput of a cluster for a given model placement. In this section, we introduce a mixed-integer linear programming (MILP)-based method to find a model placement that maximizes the max flow, thus maximizing serving throughput. The MILP formulation has a linear number of variables and constraints with respect to the number of compute nodes and network connections. The key challenges addressed include (1) formulation of system-level constraints as linear number of conditions to satisfy, (2) expression of these conditions with linear number of variables, and (3) linearization of each condition using auxiliary variables, specifically, each constraint is expressed as at most three linear constraints with the help of at most two auxiliary variables. An overview of the variables and constraints is shown in Table 5 and 6.
@@ 4.4. Optimal Model Placement with MILP
To represent the model placement on each compute node, we introduce two groups of variables in our MILP formulation. Suppose that the model has a total of $L$ layers and each compute node holds a continuous subset of the model. For each compute node $c_{i}$, we introduce an integer variable $s_{i}$ to represent the first layer $c_{i}$ holds. Suppose compute node $c_{i}$ can hold at most $k$ layers on its GPU, we further introduce $k$ binary variables $b_{i}^{1},b_{i}^{2},...,b_{i}^{k}$ to indicate the number of layers node $c_{i}$ holds ($b_{i}^{j}=1$ if $c_{i}$ holds $j$ layers). We choose to express model placement with $k$ binary variables (instead of one integer for the number of layers) because this formulation facilitates the expression of inference throughput constraints as discussed below. The end layer index of $c_{i}$ can be expressed as $e_{i}=s_{i}+\sum_{j=1}^{k}j\cdot b_{i}^{j}$. Therefore, $c_{i}$ holds layers in range $[s_{i},e_{i})$.
Figure 3. Helix overview. In Helix, the coordinator plans model placement as described in Sec. 4.4. We only need to run model placement once for each cluster. When a new request arrives, the coordinator node runs Helix scheduler to assign it a per-request pipeline and sends it to the first node in the pipeline. Each compute node in the pipeline performs inference on the request on the layers it is responsible for and sends the (output for the) request to the next node in the pipeline. When the last node in the pipeline finishes performing inference on its layers, it will send the output token for the request to the coordinator (Worker Finished). The coordinator schedules generation of the next token for the request using the same pipeline.
[…]
Figure for system overview.
@@
We introduce two groups of variables to constrain the number of inference requests that can go through each network connection. For network connection between compute node $c_{i}$ and $c_{j}$, we introduce a real variable $f_{i,j}$ to denote the amount of flow from $c_{i}^{out}$ to $c_{j}^{in}$ in the graph abstraction. We further introduce a binary variable $d_{i,j}$ to denote whether the network connection is valid (as defined in Sec. 4.3). The constraints we introduce below will use $d_{i,j}$ to ensure that requests can only be transmitted through valid connections. For network connections between coordinator node and compute nodes, we similarly introduce two variables similar to above, but replace $i$/$j$ with source/sink.
@@
To ensure that the model placement found by the MILP solver is valid, we need the following two constraints for each compute node $c_{i}$. First, $c_{i}$ should have only one valid model placement, meaning that $\sum_{j=1}^{k}b_{i}^{j}=1$. Moreover, the first and last layer $c_{i}$ holds must be within the range of $L$ layers, meaning that $0\leq s_{i}<L$ and $e_{i}\leq L$. ($c_{i}$ holds layers in range $[s_{i},e_{i})$)
@@
For each compute node $c_{i}$, the sum of flow that goes in to $c_{i}^{in}$ must be equal to that goes out of $c_{i}^{out}$ because of flow conservation. This constraint can be expressed as $\sum_{u}f_{u,i}=\sum_{v}f_{i,v}$, where $u$ and $v$ enumerate through all nodes except $i$.
@@
For compute node $c_{i}$, the amount of flow that passes through $(c_{i}^{in},c_{i}^{out})$ should not exceed its maximum inference throughput. We can impose this constraint with $\sum_{u}f_{ui}\leq\sum_{j=1}^{k}b_{i}^{j}\cdot T_{j}$. Here, $T_{j}$ is a constant that represents the maximum number of tokens node $c_{i}$ can process in one second when holding $j$ layers, which is obtained through a one-time profiling process.
@@
We need to determine the validity of network connections to know if requests can be transmitted through them. For a network connection from the coordinator node to the compute node $c_{i}$, it is valid only if $c_{i}$ holds the first layer of the model. To express this constraint with MILP, we need to linearize it into the following form: $s_{i}\leq L\cdot(1-d_{source,i})$. Similarly, for network connection from compute node $c_{i}$ to coordinator, we constrain its validity with $L\cdot d_{i,sink}\leq e_{i}$. For network connection from compute node $c_{i}$ to $c_{j}$, its validity $d_{i,j}$ is determined by whether $s_{j}\leq e_{i}<e_{j}$ holds. To linearize this condition, we need to introduce two binary auxiliary variables $cond^{1}_{i,j}$ and $cond^{2}_{i,j}$. $cond^{1}_{i,j}$ takes value 1 only if $s_{j}\leq e_{i}$, which can be linearized as $(L+1)(1-cond^{1}_{i,j})\geq s_{j}-e_{i}$. $cond^{2}_{i,j}$ takes value 1 only if $e_{i}<e_{j}$, which can be linearized as $e_{j}-e_{i}\geq 1-(L+1)(1-cond^{2}_{i,j})$. The network connection is valid only if both binary auxiliary variables are true, which can be expressed as $d_{i,j}\leq 0.5*cond^{1}_{i,j}+0.5*cond^{2}_{i,j}$.
We remark that if $s_{j}<e_{i}<e_{j}$, then requests coming from $c_{i}$ will only infer layers $[e_{i},e_{j})$ on $c_{j}$. We call this partial inference. If partial inference is not allowed, then the connection validity constraints can be simplified to $d_{i,j}=1$ only if $e_{i}=s_{j}$, which linearizes to two constraints $L\cdot d_{i,j}\leq L+s_{j}-e_{i}$ and $L\cdot d_{i,j}\leq L-s_{j}+e_{i}$.
@@
We only allow flow to pass through valid network connections, and the flow should not be larger than the connection’s maximum transmission throughput. To enforce this constraint, we add $f_{i,j}\leq d_{i,j}\cdot S_{i,j}$ as a constraint into the MILP problem. $S_{i,j}$ is the maximum number of tokens that can be transmitted through the network connection, which can be calculated via profiling and using methods mentioned in Sec. 4.3.
@@
4.5. Analyzing and Speeding up MILP